• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Singh

    Log entry time 08:51:47 on May22,2003

    Entry number 103273

    keyword=owl shift summary

    0000 target at 0 deg, in 12C Multi Foils position
    0227 beam back
    0230 2246
    0245 2247, stopped this run b/c of beam current/rate anomalies. see 103258
    0300 it appears that MCC is doing HARP scans, beam not good for data taking
    0440 septum quenched see parts 1, 2, 3, 4, 5.
    0720 changing to kin 5.15c (-2%), see 103269

    owl shift spetrum spit chart

    run time kin tar raster events
    2246 0230:0245 5.15.b.ii 12C Multi Foils Off 2.0M
    2247 0245:0256 5.15.b.ii 12C Multi Foils Off 1.1M